# Taxonomy of periodic orbits and gravitational waves in a deformed Schwarzschild black hole spacetime

### v3

@@ 1 Introduction
Since the successful detection of the gravitational-wave signal GW150914 in 2016 by LIGO and Virgo, gravitational waves have become a new messenger of cosmic explorationAbbott and others (2016b, c, a, d). These oscillations of spacetime were predicted by Einstein’s theory of general relativity and have significantly broadened the research horizon on the physics frontierBarack and others (2019); Berti and others (2015). Among all the sources for space-based gravitational wave detectors——such as TaijiHu and Wu (2017), TianqinLuo and others (2016); Gong et al. (2021), and LISASchutz (1999); Amaro-Seoane et al. (2017); Gair et al. (2004)——the Extreme Mass Ratio Inspiral system (EMRI) is a crucial class of gravitational wave sources in the mHz frequency band, in which a stellar-mass object inspirals around a supermassive black holeHughes (2001).
@@
To characterize EMRI orbits, the study of particle trajectories beyond circular orbits around black holes is of great significanceCardoso et al. (2021). It helps to understand the dynamics around black holes, providing a powerful theoretical framework for addressing challenges in astrodynamics. Among all the particle trajectories, periodic orbits are particularly important, as all eccentric orbits will undergo approximately periodic orbits during their inspiral towards the central black holeLevin and Perez-Giz (2008); Levin (2009). Periodic orbits exhibit topological characteristics of so-called ”multi-leaf clovers” which means whirling around the central black hole and zooming out in quasi elliptical orbits Barack and Cutler (2004); Glampedakis and Kennefick (2002). Three integers index orbits topologically. They represent scaling, rotation and vertex behavior, respectivelyLevin and Perez-Giz (2008). Therefore, an orbital taxonomy can associate periodic orbits with rational numbers. Using this kind of classification, a series of discussions have already been carried out in different spacetimes, such as Schwarzschild and KerrLevin and Perez-Giz (2008); Levin (2009); Compère et al. (2022), Reissner-NordströmMisra and Levin (2010), quantum-corrected black holeGuo et al. (2025); Yang et al. (2025), Einstein-Æther black holeLu and Zhu (2025), and so onLiu et al. (2019); Wei et al. (2019); Gao and Deng (2020); Lin and Deng (2022); Wang et al. (2022); Lin and Deng (2023); Tu et al. (2023); Jiang et al. (2024); Zahra et al. (2025); Deng et al. (2025); Lin and Deng (2021); Deng (2020); Melis et al. (2025); Gong et al. (2025); Ahmed et al. (2025a); Sharipov et al. (2025); Alloqulov et al. (2025). An effective way to display these orbits is to calculate gravitational waveforms, and in some cases waveforms have been studied before, see Refs.Lu and Zhu (2025); Yang et al. (2025); Zahra et al. (2025); Deng et al. (2025); Gong et al. (2025); Ahmed et al. (2025a, b) and references therein.
@@
Periodic orbits not only play an important role in orbit taxonomy but also have significance in studying transient resonances. As mentioned before, generic bound orbits can be approximately described by periodic orbits with perturbationLevin and Perez-Giz (2008). In Kerr spacetime, when radiation reaction is considered, small compact objects will go through low order periodic orbits and transient resonances as a post adiabatic effect will emerge, accompanied by mutations of the motion constants Flanagan and Hinderer (2012); Berry et al. (2016). Moreover, resonances are focused more often in non-Kerr spacetime, where the Carter symmetry is broken and the spacetime becomes non-integrable. In this case, the resonance islands will significantly generate within the phase spacePan et al. (2023).
@@
Among various studies on non-Kerr spacetime, a metric, developed by Destounis, Suvorov and KokkotasDestounis et al. (2020) (DSK), includes generic deviations from the Kerr spacetime with a parameter $\alpha_{Q}$ controlling the Carter symmetry. In their work, gravitational-wave glitches associated with $r$-$\theta$ resonant orbits have been studiedDestounis et al. (2021, 2020). However, a systematic classification of periodic orbits in this spacetime remains to be studied. Given the complexity of non-Kerr spacetime, comprehensive analysis of periodic orbits in the generic case of this kind of spacetime is a big challenge. Therefore, we explore the $r$-$\varphi$ periodic orbits of this spacetime in static limit, namely a deformed Schwarzschild spacetime. Under this circumstance, the null circular orbit can be analytically solved. Actually, we solve null and time-like circular orbits in this spacetime and find that the circular orbits will be qualitatively different from those in the Schwarzschild spacetime in some situation. We also give a relatively complete taxonomy of periodic orbits with a triplet $(z,w,v)$ describing the number of orbital branches, the orbits whirling and the movement of the small object between apastrons, respectively. Based on this taxonomy, we study how the deformation affects the shape of them. At last, we also calculate the gravitational waveform signals generated by different periodic orbits, analyze the effects of deformation on them and show the differences from Schwarzschild spacetime by calculating the mismatch, which can be used to test the DSK model in future space-based observation.
@@
The paper is constructed as follows. In section 2, a brief introduction of the deformed Schwarzschild spacetime is presented. After this, in section 3 the details of different circular orbits are analyzed separately. Firstly, the equations of motion are derived for the test particle and the effective potential from the Lagrangian is written. Then we investigate how the deformation affects the effective potential and study the marginally bound orbits (MBOs) and the marginally stable circular orbits (MSCOs). Then, in section 4 the rational number $q$ of a periodic orbit, related to the triplet, is calculated and the impact of deformation parameter on the energy and orbital angular momentum is explored. After this, a catalog of periodic orbits is concluded. In section 5 we present the calculations of gravitational waveforms and comparison with Schwarzschild spacetime scenario, illustrating the influence of modification in spacetime. Finally, the conclusions and discussions are given in section 6.
2 The deformed Schwarzschild spacetime
@@
with a parameter $\alpha$ describing the degree of deformation, which is the parameter $\alpha_{Q}$ in the non-Kerr spacetimeDestounis et al. (2020). Note that the event horizon is fixed at $r=2M$, independent of $\alpha$. This spacetime can be taken as a form of an exact, vacuum solution of mixed scalar-$f(R)$ theories Suvorov (2021).
The intuition about the physical essence of the deformation can be illustrated by comparison with existing solutions. The $tt$-component is the limit of a particular non-vacuum black hole metric introduced by Bardeen, of which the $tt$-component reads $g_{tt}=-1+2M/r-3Mq^{2}/r^{3}+\mathcal{O}(1/r^{4})$. The Bardeen solution involves a metric coupled to a nonlinear electromagnetic field which is generated by Bardeen magnetic monopole. Therefore, the parameter $\alpha$ can be explained as a gravitational analogy of a Bardeen magnetic monopoleDestounis et al. (2020). More detailed analysis needs tools like multipole moment expansionsThorne (1980).
@@ 3.1 The Effective Potential
3.2 Different Circular Orbits
@@ 4 Periodic Orbits
5 Gravitational Radiation from periodic orbits
@@
The numerical Kludge model will be introduced briefly primarily, since it captures many qualitative features of EMRIsChua et al. (2017); Babak et al. (2007). It can even be available in self-force condition by including the particle dynamics. Although the self-consistency will degrade, it is still effective as an approximation to Teukolsky-based waveform in strong fieldChua et al. (2017). To compute the waveform, equations of motion (4) need be solved numerically to construct trajectory in space of motion constants firstly. As mentioned before, the influence caused by radiation reaction can be ignored. Secondly, the curved-space coordinates of trajectory then will be associated artificially with coordinates in flat space. Using the standard quadrupole formula, the corresponding gravitational waveform is generated.
The symmetric trace-free (STF) mass quadrupole of the small object with mass $m$ isThorne (1980):
@@
where $\iota$ and $\eta$ represent the inclination angle and the longitude of the perihelion of the orbital plane, respectivelyBarack and Cutler (2004). Then the gravitational wave polarizations can be obtained as follows:
@@
Considering the transverse-traceless gauge, the concrete expressions of crucial components are given asBabak et al. (2007):
@@
In order to show the influence of the deformation parameter on the gravitational waveforms of different periodic orbits, we consider a small compact object with mass $m=10M_{\odot}$ around a supermassive black hole with mass $M=10^{6}M_{\odot}$ at the luminosity distance $D_{\text{L}}=1\text{Gpc}$. Take a simple case, the inclination angle $\iota$ is set as $\pi/2$ and the longitude of perihelion is fixed at 0. Moreover, the polarization angle is also set as 0 for simplicityBabak et al. (2007). To show contrast, we fix the energy at $E=\sqrt{0.9755751102284231}$ and change the deviation parameter $\alpha$. When $\alpha$ is large enough, the perihelion is too small, as shown in Fig.6. So numerical kludge model will not work. And then the waveforms are represented by two polarization components $h_{+}$ and $h_{\times}$ to show their characteristics in Fig.7. We list three parameter conditions and their waveform contrast here.
@@
In Figs 6 and 7, a typical kind of periodic orbits is plotted with the triplet $(1,2,0)$ and their $h_{+}$ and $h_{\times}$ modes with different deformation parameters. It shows that a clear relationship exists between gravitational waveforms and orbit. The giant leaf results in very long smooth phases. And there are rapid oscillations which align with the zoom and whirl behaviors. Moreover, it is clear that in comparison with the Schwarzschild case the parameter influences the phase of the waveforms mainly and the waveform amplitudes in a slight way when parameter increases. To clearly quantitatively illustrate the differences in gravitational waveforms with different parameters, a useful characteristic is the mismatch between two waveformsAllen (2019); Qiao et al. (2025); Lai et al. (2025). The specific calculation formula is given asPompili (2025):
@@
Although not shown in Fig. 8 , we find that the mismatch exceeds $0.05$ for $\alpha\geq 0.4$. The mismatch increases with the deformation parameter, as expected for progressively larger departures from the Schwarzschild situation. We also find that the mismatch varies across different periodic orbits, indicating an orbit-dependent sensitivity to the deformation. It is noteworthy that with the same number of $w$ and $v$, the mismatch of the orbit with three branches is greater than that of the orbit with two branches, which may be attributed to the accumulation of waveform differences induced by the number of orbital branches. While the deformation induces only modest changes in the waveform amplitude and phase, the mismatch provides a preliminary means of discriminating among these cases. A definitive assessment, however, requires a dedicated analysis of detectability and measurability, following and extending the methodology of Ref.Speeney et al. (2022). Overall, these results suggest a plausible pathway toward probing subtle physical features in future gravitational wave detection.
6 Conclusions and Discussions
@@
The work presented here reveals some distinct features of geodesics from the Schwarzschild black hole. Nevertheless, there are some intriguing problems to address in the future. For example, orbital resonances can be taken into further investigationFlanagan and Hinderer (2012). Moreover, the possible mechanism in the inner region and the disappearance of the photon ring are also worth considering. Last but not the least, while research for geodesics in DSK metric under general conditions is challenging, qualitative analysis can be taken into consideration. We leave these topics to future work.
7 Acknowledgements

### v4

@@ I Introduction
Since the successful detection of the gravitational-wave signal GW150914 in 2016 by LIGO and Virgo, gravitational waves have become a new messenger of cosmic exploration[2, 3, 1, 4]. These oscillations of spacetime were predicted by Einstein’s theory of general relativity and have significantly broadened the research horizon on the physics frontier[12, 14]. Among all the sources for space-based gravitational wave detectors——such as Taiji[29], Tianqin[40, 27], and LISA[46, 9, 23]——the Extreme Mass Ratio Inspiral system (EMRI) is a crucial class of gravitational wave sources in the mHz frequency band, in which a stellar-mass object inspirals around a supermassive black hole[30].
@@
To characterize EMRI orbits, the study of particle trajectories beyond circular orbits around black holes is of great significance[15]. It helps to understand the dynamics around black holes, providing a powerful theoretical framework for addressing challenges in astrodynamics. Among all the particle trajectories, periodic orbits are particularly important, as all eccentric orbits will undergo approximately periodic orbits during their inspiral towards the central black hole[33, 34]. Periodic orbits exhibit topological characteristics of so-called ”multi-leaf clovers” which means whirling around the central black hole and zooming out in quasi elliptical orbits [11, 25]. Three integers index orbits topologically. They represent scaling, rotation and vertex behavior, respectively[33]. Therefore, an orbital taxonomy can associate periodic orbits with rational numbers. Using this kind of classification, a series of discussions have already been carried out in different spacetimes, such as Schwarzschild and Kerr[33, 34, 17], Reissner-Nordström[42], quantum-corrected black hole[28, 54], Einstein-Æther black hole[39], and so on[38, 53, 24, 36, 52, 37, 51, 31, 55, 18, 35, 19, 41, 26, 5, 47, 8]. An effective way to display these orbits is to calculate gravitational waveforms, and in some cases waveforms have been studied before, see Refs.[39, 54, 55, 18, 26, 5, 6] and references therein.
@@
Periodic orbits not only play an important role in orbit taxonomy but also have significance in studying transient resonances. As mentioned before, generic bound orbits can be approximately described by periodic orbits with perturbation[33]. In Kerr spacetime, when radiation reaction is considered, small compact objects will go through low order periodic orbits and transient resonances as a post adiabatic effect will emerge, accompanied by mutations of the motion constants [22, 13]. Moreover, resonances are focused more often in non-Kerr spacetime, where the Carter symmetry is broken and the spacetime becomes non-integrable. In this case, the resonance islands will significantly generate within the phase space[43].
@@
Among various studies on non-Kerr spacetime, a metric, developed by Destounis, Suvorov and Kokkotas[20] (DSK), includes generic deviations from the Kerr spacetime with a parameter $\alpha_{Q}$ controlling the Carter symmetry. In their work, gravitational-wave glitches associated with $r$-$\theta$ resonant orbits have been studied[21, 20]. However, a systematic classification of periodic orbits in this spacetime remains to be studied. Given the complexity of non-Kerr spacetime, comprehensive analysis of periodic orbits in the generic case of this kind of spacetime is a big challenge. Therefore, we explore the $r$-$\varphi$ periodic orbits of this spacetime in static limit, namely a deformed Schwarzschild spacetime. Under this circumstance, the null circular orbit can be analytically solved. Actually, we solve null and time-like circular orbits in this spacetime and find that the circular orbits will be qualitatively different from those in the Schwarzschild spacetime in some situation. We also give a relatively complete taxonomy of periodic orbits with a triplet $(z,w,v)$ describing the number of orbital branches, the orbits whirling and the movement of the small object between apastrons, respectively. Based on this taxonomy, we study how the deformation affects the shape of them. At last, we also calculate the gravitational waveform signals generated by different periodic orbits, analyze the effects of deformation on them and show the differences from Schwarzschild spacetime by calculating the mismatch, which can be used to test the DSK model in future space-based observation.
@@
The paper is constructed as follows. In section II, a brief introduction of the deformed Schwarzschild spacetime is presented. After this, in section III the details of different circular orbits are analyzed separately. Firstly, the equations of motion are derived for the test particle and the effective potential from the Lagrangian is written. Then we investigate how the deformation affects the effective potential and study the marginally bound orbits (MBOs) and the marginally stable circular orbits (MSCOs). Then, in section IV the rational number $q$ of a periodic orbit, related to the triplet, is calculated and the impact of deformation parameter on the energy and orbital angular momentum is explored. After this, a catalog of periodic orbits is concluded. In section V we present the calculations of gravitational waveforms and comparison with Schwarzschild spacetime scenario, illustrating the influence of modification in spacetime. Finally, the conclusions and discussions are given in section VI.
II The deformed Schwarzschild spacetime
@@
with a parameter $\alpha$ describing the degree of deformation, which is the parameter $\alpha_{Q}$ in the non-Kerr spacetime[20]. Note that the event horizon is fixed at $r=2M$, independent of $\alpha$. This spacetime can be taken as a form of an exact, vacuum solution of mixed scalar-$f(R)$ theories [49].
The intuition about the physical essence of the deformation can be illustrated by comparison with existing solutions. The $tt$-component is the limit of a particular non-vacuum black hole metric introduced by Bardeen, of which the $tt$-component reads $g_{tt}=-1+2M/r-3Mq^{2}/r^{3}+\mathcal{O}(1/r^{4})$. The Bardeen solution involves a metric coupled to a nonlinear electromagnetic field which is generated by Bardeen magnetic monopole. Therefore, the parameter $\alpha$ can be explained as a gravitational analogy of a Bardeen magnetic monopole[20]. More detailed analysis needs tools like multipole moment expansions[50].
@@ III.1 The Effective Potential
III.2 Different Circular Orbits
@@ IV Periodic Orbits
V Gravitational Radiation from periodic orbits
@@
The numerical Kludge model will be introduced briefly primarily, since it captures many qualitative features of EMRIs[16, 10]. It can even be available in self-force condition by including the particle dynamics. Although the self-consistency will degrade, it is still effective as an approximation to Teukolsky-based waveform in strong field[16]. To compute the waveform, equations of motion (4) need be solved numerically to construct trajectory in space of motion constants firstly. As mentioned before, the influence caused by radiation reaction can be ignored. Secondly, the curved-space coordinates of trajectory then will be associated artificially with coordinates in flat space. Using the standard quadrupole formula, the corresponding gravitational waveform is generated.
The symmetric trace-free (STF) mass quadrupole of the small object with mass $m$ is[50]:
@@
where $\iota$ and $\eta$ represent the inclination angle and the longitude of the perihelion of the orbital plane, respectively[11]. Then the gravitational wave polarizations can be obtained as follows:
@@
Considering the transverse-traceless gauge, the concrete expressions of crucial components are given as[10]:
@@
In order to show the influence of the deformation parameter on the gravitational waveforms of different periodic orbits, we consider a small compact object with mass $m=10M_{\odot}$ around a supermassive black hole with mass $M=10^{6}M_{\odot}$ at the luminosity distance $D_{\text{L}}=1\text{Gpc}$. Take a simple case, the inclination angle $\iota$ is set as $\pi/2$ and the longitude of perihelion is fixed at 0. Moreover, the polarization angle is also set as 0 for simplicity[10]. To show contrast, we fix the energy at $E=\sqrt{0.9755751102284231}$ and change the deviation parameter $\alpha$. When $\alpha$ is large enough, the perihelion is too small, as shown in Fig.6. So numerical kludge model will not work. And then the waveforms are represented by two polarization components $h_{+}$ and $h_{\times}$ to show their characteristics in Fig.7. We list three parameter conditions and their waveform contrast here.
@@
In Figs 6 and 7, a typical kind of periodic orbits is plotted with the triplet $(1,2,0)$ and their $h_{+}$ and $h_{\times}$ modes with different deformation parameters. It shows that a clear relationship exists between gravitational waveforms and orbit. The giant leaf results in very long smooth phases. And there are rapid oscillations which align with the zoom and whirl behaviors. Moreover, it is clear that in comparison with the Schwarzschild case the parameter influences the phase of the waveforms mainly and the waveform amplitudes in a slight way when parameter increases. To clearly quantitatively illustrate the differences in gravitational waveforms with different parameters, a useful characteristic is the mismatch between two waveforms[7, 45, 32]. The specific calculation formula is given as[44]:
@@
Although not shown in Fig. 8 , we find that the mismatch exceeds $0.05$ for $\alpha\geq 0.4$. The mismatch increases with the deformation parameter, as expected for progressively larger departures from the Schwarzschild situation. We also find that the mismatch varies across different periodic orbits, indicating an orbit-dependent sensitivity to the deformation. It is noteworthy that with the same number of $w$ and $v$, the mismatch of the orbit with three branches is greater than that of the orbit with two branches, which may be attributed to the accumulation of waveform differences induced by the number of orbital branches. While the deformation induces only modest changes in the waveform amplitude and phase, the mismatch provides a preliminary means of discriminating among these cases. A definitive assessment, however, requires a dedicated analysis of detectability and measurability, following and extending the methodology of Ref.[48]. Overall, these results suggest a plausible pathway toward probing subtle physical features in future gravitational wave detection.
VI Conclusions and Discussions
@@
The work presented here reveals some distinct features of geodesics from the Schwarzschild black hole. Nevertheless, there are some intriguing problems to address in the future. For example, orbital resonances can be taken into further investigation[22]. Moreover, the possible mechanism in the inner region and the disappearance of the photon ring are also worth considering. Last but not the least, while research for geodesics in DSK metric under general conditions is challenging, qualitative analysis can be taken into consideration. We leave these topics to future work.
Acknowledgements